# Decomposing Representation Space into Interpretable Subspaces with Unsupervised Learning

### v1

@@ Abstract
Understanding internal representations of neural models is a core interest of mechanistic interpretability. Due to its large dimensionality, the representation space can encode various aspects about inputs. To what extent are different aspects organized and encoded in separate subspaces? Is it possible to find these “natural” subspaces in a purely unsupervised way? Somewhat surprisingly, we can indeed achieve this and find interpretable subspaces by a seemingly unrelated training objective. Our method, neighbor distance minimization (NDM), learns non-basis-aligned subspaces in an unsupervised manner. Qualitative analysis shows subspaces are interpretable in many cases, and encoded information in obtained subspaces tends to share the same abstract concept across different inputs, making such subspaces similar to “variables” used by the model. We also conduct quantitative experiments using known circuits in GPT-2; results show a strong connection between subspaces and circuit variables. We also provide evidence showing scalability to 2B models by finding separate subspaces mediating context and parametric knowledge routing. [^1] Code and all trained matrices are available at https://github.com/huangxt39/SubspacePartition Viewed more broadly, our findings offer a new perspective on understanding model internals and building circuits.
@@ 1 Introduction
Researchers have made significant progress in this direction in recent years, uncovering circuits consisting of various types of basic units (or mediators (Mueller et al., 2024)), including components (attention heads and MLPs) of transformers (Wang et al., 2022; Hanna et al., 2023), sparse features obtained from Sparse Auto-Encoders (SAEs) or its variants (Cunningham et al., 2023; Bricken et al., 2023; Dunefsky et al., 2024; Ameisen et al., 2025), and subspaces (Geiger et al., 2024; Wu et al., 2023; Geiger et al., 2025). While providing valuable insights, they are limited in different ways. The information transferred between components is hard to understand, and researchers need to design specialized counterfactual data targeting certain aspect of the input to perform causal intervention, or use easy but much less reliable methods such as observing attention weights or projecting to vocabulary space. Sparse features provide a direct way to interpret (by reading from the inputs they activate on), but resulting circuits are input-dependent, in other words, specific for each input sequence, as different features are activated in different inputs. This approach also usually modifies the model’s computation, replacing activations or MLPs with reconstruction from sparsity-encouraged new modules, in order to better analyze feature-to-feature interactions. Lastly, current subspace-based approaches require supervision from a abstract causal model specified by human. The abstract model represents a clear hypothesis of how the real model works, thus requiring anticipating possible mechanisms. We refer to (Mueller et al., 2024) for a good summary of different mediators.
@@
In this paper, we explore a new direction: unsupervised representation space decomposition. We show we can learn a partitioning of representation space by doing neighbor distance minimization (NDM) among subspace activations. We explain the intuition behind this training objective in terms of feature groups and mutual information and show that it looks for a partition such that subspaces are as independent as possible. Similar to (Geiger et al., 2024), an orthogonal matrix is learned, which rotates and reflects the space before partitioning on the transformed basis. It captures the distributed nature of neural representations. While they aim to find the subspace for certain target variables by supervised learning, our method is unsupervised and makes use of “natural” structure inside the model. We first demonstrate the effectiveness of NDM in a toy setting, then move on to real language models. Somewhat surprisingly, we find that there is indeed meaningful inner structure of representation space in real-world models. We show that resulting subspaces are interpretable with InversionView (Huang et al., 2024b), and we quantitatively test alignment between subspaces and variables in known circuits in GPT-2 Small using causal intervention. Moreover, we demonstrate its applicability on larger 2B models by showing that NDM finds separate subspaces mediating context knowledge and parametric knowledge.
@@ 2 Motivation and Background
$\begin{array}[]{c}\boldsymbol{h}=\mathbf{W}\boldsymbol{x}\\ \boldsymbol{x}^{\prime}=\mathrm{ReLU}(\mathbf{W}^{T}\boldsymbol{h}+\boldsymbol{b})\end{array}$ (1)
@@
In this toy setting, Elhage et al. (2022) found that sparsity is necessary to allow superposition to occur. In this paper, we argue that mutual exclusiveness could be a more fundamental condition. More specifically, for a group of features, if only one (or none) occurs at a time, they are said to be mutually exclusive, and can be represented with superposition without much loss of information. Although sparsity encourages mutual exclusiveness, they are fundamentally different.
If mutual exclusiveness holds, superposition works well; the only discrepancy between $\boldsymbol{x}$ and $\boldsymbol{x}^{\prime}$ is because of small interference between almost orthogonal features. On the contrary, if multiple features are activated, they might cancel each other in certain directions, causing loss of information, so that reconstructed $\boldsymbol{x}^{\prime}$ can be totally different from $x$ (Fig. 5(a) and 5(b) in App.). In other words, if two or more features often activate together, the strategy of superposition would not work well, and the model would be forced to represent only part of the features with orthogonal direction.
@@ 3 Methodology
Figure 1: (a): $\mathbf{W}^{T}\mathbf{W}$, $\mathbf{W}\in\mathbb{R}^{12\times 40}$ (Sec. 2). Each column of $\mathbf{W}$ is a feature direction, its first 20 columns belong to one feature group, the last 20 to the other. We can see the bottom left and top right parts are zero, which means feature vectors of different groups are orthogonal. (b): Illustration for the intuition of our method (Sec. 3), correct partition can make data points easier to model. (c) (d): $\mathbf{R}\mathbf{W}$ before and after NDM training (Sec. 4). The final dimension configuration is $d_{1}=6,d_{2}=6$. The corresponding toy model is shown in (a), whose d=12 and group=(20,20). We can see after training, the first subspace (first 6 coordinates) only depends on the first feature group (first 20 coordinates), and likewise for the other subspace.
@@
We define a partition as a tuple consisting of an orthogonal matrix $\mathbf{R}$ and a dimension configuration $c$ (number of subspaces and their dimensionality). $\mathbf{R}$ rotates and reflects the space as needed (similar to (Geiger et al., 2024)), then the transformed space is partitioned on its standard basis (simply grouping coordinates) according to $c$. We optimize $\mathbf{R}$ as follows: given $N$ model activations $[\boldsymbol{h_{1}}\cdots\boldsymbol{h_{N}}]$ where $\boldsymbol{h_{n}}\in\mathbb{R}^{d}$, and an orthogonal matrix to be trained $\mathbf{R}\in\mathbb{R}^{d\times d}$  [^2] Libraries like Pytorch (Paszke et al., 2019) contain parametrization to ensure orthogonality of matrix, dimension configuration $c=[d_{1},\cdots,d_{S}]$, where $S$ is the number of subspaces, $\sum_{s=0}^{S}d_{s}=d$, we do
$\begin{array}[]{c}[\boldsymbol{\hat{h}_{1}}\cdots\boldsymbol{\hat{h}_{n}}]=\mathbf{R}[\boldsymbol{h_{1}}\cdots\boldsymbol{h_{n}}]\end{array}$ (2)
$\hat{\boldsymbol{h}}_{n}=\begin{bmatrix}\hat{\boldsymbol{h}}_{n}^{(1)}\\ \vdots\\ \hat{\boldsymbol{h}}_{n}^{(S)}\end{bmatrix},\quad\hat{\boldsymbol{h}}_{n}^{(s)}\in\mathbb{R}^{d_{s}}$ (3)
@@
However, there is still a question: how to find the correct dimension configuration (number of subspaces and their dimensions)? It is not differentiable and is unclear what is the appropriate objective to optimize for. In our method, we use MI to address this problem. The procedure is as follows: we start with a configuration of small and equal-sized subspaces and train the orthogonal matrix. We measure the MI between subspaces regularly during training using KSG estimator (Kraskov et al., 2004). If the MI [^3] In practice, we divide the mutual information value by the sum of the two subspaces’ dimensions. is higher than a threshold, we merge the two subspaces, so the nearest neighbor should be found in the joint subspace afterwards and keep training the orthogonal matrix. We stop training when we cannot find a pair of subspaces whose MI is above the threshold. This procedure is in line with our goal of reducing the dependency between subspaces, as we merge them if their MI is unavoidably high. We describe our method in detail in Algorithm 1 and App. B. We also experimented with other methods for determining the configuration, see App. F.3. In summary, NDM finds a partition containing both $\mathbf{R}$ and $c$.
@@ 4 Experiments in Toy Models
[…]
The result for the toy model mentioned in Sec.2 is shown in Fig. 1. We initialize $\mathbf{R}$ as an identity matrix, and the $\mathbf{R}\mathbf{W}$ after training is almost perfect. We present more experiments in Fig. 9-12 in App. C.2, where we apply NDM on different toy models with various numbers of features and groups. The results overall show a strong performance of NDM, we are able to find the subspace partition close to the ground truth in all cases. In sum, we show that in the toy setting where there exist orthogonal subspaces dedicated to each group, we can find these subspaces by NDM which only relies on model activations.
@@ 5 Experiments in Language Models
The key question is NDM’s applicability to real-world neural models. Unlike the toy setting, we do not have access to ground-truth features—how do we evaluate our method? We design new ways to do evaluation and present strong evidence showing NDM finds interpretable subspaces, many of them corresponding to meaningful concepts. In this section, we denote residual stream activations as $\boldsymbol{h}^{l,\{mid,post\}}$, where $l$ is the layer index (starting from 0), $mid$ means after adding the attention sublayer’s output, and $post$ means after adding the MLP sublayer’s output.
@@ 5.1 Quantitative Evaluation Based on GPT-2 Circuits
The evaluation is based on subspace activation patching, or subspace interchange intervention (Geiger et al., 2024). The main intuition is that, when processing inputs, key intermediate results should ideally lie in a single subspace. For example, in an induction head circuit where the residual stream contains the previous token, the previous token should be placed in a single subspace, independent of which token it is from. Therefore, if we patch subspace activations with the value they take in sequences with different previous tokens, the effect should be concentrated on that correct subspace. We refer to the previous token as the target information, which is changed in the counterfactual input. Different target information should ideally be concentrated in different subspaces. We describe subspace patching in detail in App. D.1 and address potential concerns from (Makelov et al., 2023) in App. F.2. In sum, we aim to explore whether the high-level “variable” lies in the same subspace no matter what value it takes by measuring inequality or concentration of subspace patching effect.
@@
We construct a test suite, making use of known circuits in GPT2 Small. It consists of 5 tests. The first 4 tests are based on the IOI circuit (Wang et al., 2022), where we intervene on the following pieces of information respectively: (1) previous token information contained in layer 4 residual stream at the token after the first subject name, (2) subject position information contained in layer 6 residual stream at the second subject name, (3) subject position and (4) subject name information contained in layer 8 residual stream at the last token. The last test is based on Greater-than circuit (Hanna et al., 2023): (5) we change information about the last two digits of the starting year in layer 9 residual stream at the last token. We measure patching effect as the change in logit difference for IOI-based tests and change in valid answer probability for the Greater-than-based test. See more details in in App. D.2. We then quantify the inequality of patching effect using Gini coefficient. Though very unconventional in machine learning, we find Gini coefficient a suitable metric. Because we want to measure how strongly patching effect is focused on just one or a few subspaces, this notion is captured by Gini coefficient as it measures inequality or degree of concentration. If the target information is mainly encoded in one subspace, the patching effect for this one would be much higher than others, thus resulting in Gini coefficient close to 1. On the other hand, evenly distributed patching effect results in 0. In App. D.4 we show examples of patching effect distributions and their corresponding Gini coefficients, from which we conclude that the coefficient should be >0.6>0.6> 0.6 to be considered satisfactory (i.e., the effect is mainly concentrated in one subspace). We also calculate the Gini coefficient over patching effect normalized by dimension $d_{s}$ and variance $\text{Var}_{s}$ of subspaces.
@@
- $d_{s}$ $\text{Var}_{s}$ $d_{s}$ $\text{Var}_{s}$ $d_{s}$ $\text{Var}_{s}$ $d_{s}$ $\text{Var}_{s}$ $d_{s}$ $\text{Var}_{s}$
@@
Table 1: Results on GPT-2 Small test suite, higher is better. Test 1-5 are measuring the concentration of different target information in various places in GPT-2 Small (details in App. D.2) based on ground truth circuits (Wang et al., 2022; Hanna et al., 2023). Numbers are Gini coefficient over subspace patching effects (“-”), over effects normalized by subspace dimension (“$d_{s}$”), and over effects normalized by subspace variance (“$\text{Var}_{s}$”). Last column is the average values over the whole row. We compare NDM with 4 baselines. An identity matrix (“Identity”); a random orthogonal matrix (“Random”); PCA 1 and 2 are matrices composed of eigenvectors after applying PCA on activations. We use dimension configuration $[d_{1},\dots,d_{S}]$ of NDM result as the configuration for all baselines. $[d_{1},\dots,d_{S}]$ is sorted in descending order, eigenvectors can be sorted in either way according to eigenvalues, resulting in PCA 1 and 2 – meaning either that the smallest subspace receives the biggest (“PCA 1”) or smallest (“PCA 2”) directions of variance. Importantly, as shown in App. D.4, the degree of concentration of target information becomes satisfactory when Gini coefficient >0.6>0.6> 0.6.

### v2

@@ Abstract
Understanding internal representations of neural models is a core interest of mechanistic interpretability. Due to its large dimensionality, the representation space can encode various aspects about inputs. To what extent are different aspects organized and encoded in separate subspaces? Is it possible to find these “natural” subspaces in a purely unsupervised way? Somewhat surprisingly, we can indeed achieve this and find interpretable subspaces by a seemingly unrelated training objective. Our method, neighbor distance minimization (NDM), learns non-basis-aligned subspaces in an unsupervised manner. Qualitative analysis shows subspaces are interpretable in many cases, and encoded information in obtained subspaces tends to share the same abstract concept across different inputs, making such subspaces similar to “variables” used by the model. We also conduct quantitative experiments using known circuits in GPT-2; results show a strong connection between subspaces and circuit variables. We also provide evidence showing scalability to 2B models by finding separate subspaces mediating context and parametric knowledge routing. [^1] Link for code and trained matrices https://github.com/huangxt39/SubspacePartition Viewed more broadly, our findings offer a new perspective on understanding model internals and building circuits.
@@ 1 Introduction
Researchers have made significant progress in this direction in recent years, uncovering circuits consisting of various types of basic units (or mediators (Mueller et al., 2024)), including components (attention heads and MLPs) of transformers (Wang et al., 2022; Hanna et al., 2023), sparse features obtained from Sparse Auto-Encoders (SAEs) or its variants (Cunningham et al., 2023; Bricken et al., 2023; Dunefsky et al., 2024; Ameisen et al., 2025), and subspaces (Geiger et al., 2024; Wu et al., 2023; Geiger et al., 2025). While providing valuable insights, they are limited in different ways. The information transferred between components is hard to understand, and researchers need to design specialized counterfactual data targeting certain aspect of the input to perform causal intervention, or use easy but much less reliable methods such as observing attention weights or projecting to vocabulary space. Sparse features provide a direct way to interpret (by reading from the inputs they activate on), but resulting circuits are input-dependent, in other words, specific for each input sequence, as different features are activated in different inputs. This approach also usually modifies the model’s computation, replacing activations or MLPs with reconstruction from sparsity-encouraged new modules, in order to better analyze feature-to-feature interactions. Lastly, current subspace-based approaches require supervision from an abstract causal model specified by human. The abstract model represents a clear hypothesis of how the real model works, thus requiring anticipating possible mechanisms. We refer to Mueller et al. (2024) for a good summary of different mediators.
@@
In this paper, we explore a new direction: unsupervised representation space decomposition. We show we can learn a partitioning of representation space by doing neighbor distance minimization (NDM) among subspace activations. We explain the intuition behind this training objective in terms of feature groups and mutual information and show that it looks for a partition such that subspaces are as independent as possible. Similar to Geiger et al. (2024), an orthogonal matrix is learned, which rotates and reflects the space before partitioning on the transformed basis. It captures the distributed nature of neural representations. While they aim to find the subspace for certain target variables by supervised learning, our method is unsupervised and makes use of “natural” structure inside the model. We first demonstrate the effectiveness of NDM in a toy setting, then move on to real language models. Somewhat surprisingly, we find that there is indeed meaningful inner structure of representation space in real-world models. We show that resulting subspaces are usually interpretable with InversionView (Huang et al., 2024b). We quantitatively test alignment between subspaces and variables in known circuits in GPT-2 Small using causal intervention. Moreover, we demonstrate its applicability on larger 2B models by showing that NDM finds separate subspaces mediating context knowledge and parametric knowledge.
@@ 2 Motivation and Background
$\boldsymbol{h}=$ $\mathbf{W}\boldsymbol{x}$ $\boldsymbol{x}^{\prime}=\mathrm{ReLU}(\mathbf{W}^{T}\boldsymbol{h}+\boldsymbol{b})$ (1)
@@
In this toy setting, (Elhage et al., 2022) found that sparsity is necessary to allow superposition to occur. In this paper, we argue that mutual exclusiveness could be a more fundamental condition. More specifically, for a group of features, if only one (or none) occurs at a time, they are said to be mutually exclusive, and can be represented with superposition without much loss of information. Although sparsity encourages mutual exclusiveness, they are different, since the former concept treats features uniformly while the latter emphasizes a non-uniform view under which some sets of features can be more dependent on each other.
If mutual exclusiveness holds, superposition works well; the only discrepancy between $\boldsymbol{x}$ and $\boldsymbol{x}^{\prime}$ is because of small interference between almost orthogonal features. On the contrary, if multiple features are activated, they might cancel each other in certain directions, causing loss of information, so that reconstructed $\boldsymbol{x}^{\prime}$ can be totally different from $x$ (Fig. 6(a) and 6(b) in App.). In other words, given two or more specific features, if they often co-occur, superposing these features would not work well – incentivizing the model to represent them more orthogonally. On the other hand, if they are mutually exclusive (because the concepts they represent are mutually exclusive in the real world), the model would tend to make use of this fact and superpose their features.
@@ 3 Methodology
Figure 1: (a): $\mathbf{W}^{T}\mathbf{W}$, $\mathbf{W}\in\mathbb{R}^{12\times 40}$ (Sec. 2). Each column of $\mathbf{W}$ is a feature direction, its first 20 columns belong to one feature group, the last 20 to the other. We can see the bottom left and top right parts are zero, which means feature vectors of different groups are orthogonal. (b): Illustration for the intuition of our method (Sec. 3), correct partition can make data points easier to model. (c) (d): $\mathbf{R}\mathbf{W}$ before and after NDM training (Sec. 4). The final dimension configuration is $d_{1}=6,d_{2}=6$. The corresponding toy model is shown in (a). After training, the first subspace (first 6 coordinates) only depends on the first feature group (first 20 coordinates), and likewise for the other subspace.
@@
We define a partition as a tuple consisting of an orthogonal matrix $\mathbf{R}$ and a dimension configuration $c$ (number of subspaces and their dimensionality). $\mathbf{R}$ rotates and reflects the space as needed (similar to (Geiger et al., 2024)), then the transformed space is partitioned on its standard basis (simply grouping coordinates) according to $c$. We optimize $\mathbf{R}$ as follows: given $N$ model activations $\boldsymbol{h_{1}}\cdots\boldsymbol{h_{N}}]$ where $\boldsymbol{h_{n}}\in\mathbb{R}^{d}$, and an orthogonal matrix to be trained $\mathbf{R}\in\mathbb{R}^{d\times d}$  [^2] Libraries like Pytorch (Paszke et al., 2019) contain parametrization to ensure orthogonality of matrix, dimension configuration $c=[d_{1},\cdots,d_{S}]$, where $S$ is the number of subspaces, $\sum_{s=0}^{S}d_{s}=d$, we do
$\boldsymbol{\hat{h}_{1}}\cdots\boldsymbol{\hat{h}_{n}}]=\mathbf{R}[\boldsymbol{h_{1}}\cdots\boldsymbol{h_{n}}],\qquad\hat{\boldsymbol{h}}_{n}=\begin{bmatrix}\hat{\boldsymbol{h}}_{n}^{(1)}\\ \vdots\\ \hat{\boldsymbol{h}}_{n}^{(S)}\end{bmatrix},\quad\hat{\boldsymbol{h}}_{n}^{(s)}\in\mathbb{R}^{d_{s}}$ (2)
@@
However, there is still a question: how to find the correct dimension configuration (number of subspaces and their dimensions)? It is not differentiable and is unclear what is the appropriate objective to optimize for. In our method, we use MI to address this problem. The procedure is as follows: we start with a configuration of small and equal-sized subspaces and train the orthogonal matrix. We measure the MI between subspaces regularly during training using KSG estimator (Kraskov et al., 2004). If the MI [^3] In practice, we divide the mutual information value by the sum of the two subspaces’ dimensions. is higher than a threshold, we merge the two subspaces, so the nearest neighbor should be found in the joint subspace afterwards and keep training the orthogonal matrix. We stop training when we cannot find a pair of subspaces whose MI is above the threshold. This procedure is in line with our goal of reducing the dependency between subspaces, as we merge them if their MI is unavoidably high. We describe our method in detail in Algorithm 1 and App. B. We also experimented with other methods for determining the configuration, see App. G.4. In sum, NDM finds a partition containing both $\mathbf{R}$ and $c$.
@@ 4 Experiments in Toy Models
We first test our method in the toy model setting described in Sec.2, using a slightly different and less scalable version of Algorithm 1 (see details in App. C.2). We use Euclidean distance as the $\text{dist}(\cdot)$ in Eq. 3 and 4, because we find cosine similarity to perform worse in preliminary experiments. The toy model setting is an ideal testbed as we have access to the ground truth feature vectors and subspaces. As described, we rotate and reflect the hidden activations $\boldsymbol{h}$ from Eq.1, which results in $\boldsymbol{\hat{h}}=\mathbf{R}\mathbf{W}\boldsymbol{x}$. Subspace activations, which are defined to be a continuous span of coordinates of $\boldsymbol{\hat{h}}$, should ideally result from only one feature group. Therefore, a good $\mathbf{R}$ means rows in $\mathbf{R}\mathbf{W}$ corresponding to one subspace (subvectors of $\boldsymbol{\hat{h}}$) have nonzero values only at columns corresponding to one feature group (subvectors of $\boldsymbol{x}$). So, the important question is, can we find subspaces encoding each feature group by minimizing distances? Despite our argument before, this still sounds uncertain a priori, as we do not optimize $\mathbf{R}$ to associate subspaces with underlying groups but rather optimize the indirect objective of neighbor distance, as we want to avoid using $\mathbf{W}$ and $\boldsymbol{x}$. Somewhat beyond our expectation, the answer is yes, we find good subspaces by using NDM.
The result is shown in Fig. 1. We initialize $\mathbf{R}$ as an identity matrix, and the $\mathbf{R}\mathbf{W}$ after training is almost perfect. We present more experiments in Fig. 10-13 in App. C.2, where numbers of features and groups vary. The results overall show a strong performance of NDM, we are able to find the subspace partition close to the ground truth in all cases. In sum, in the toy setting where there exist orthogonal subspaces dedicated to each group, we indeed can find these subspaces by NDM.
@@ 5 Experiments in Language Models
The key question is NDM’s applicability to real-world neural models. Unlike the toy setting, we do not know ground-truth features, so we design new ways to do evaluation. In this section, we denote residual stream activations as $\boldsymbol{h}^{l,\{mid,post\}}$, where $l$ is the layer index (starting from 0), midmid means after adding the attention sublayer’s output, and postpost means after adding the MLP sublayer’s output.
@@ 5.1 Quantitative Evaluation Based on GPT-2 Circuits
The evaluation is based on subspace activation patching, or subspace interchange intervention (Geiger et al., 2024). The main intuition is that, when processing inputs, key intermediate results should ideally lie in a single subspace. For example, in an induction head circuit where the residual stream contains the previous token, the previous token should be placed in a single subspace, independent of which token it is from. Therefore, if we patch subspace activations with the value they take in sequences with different previous tokens, the effect should be concentrated on that correct subspace. We refer to the previous token as the target information, which is changed in the counterfactual input. Different target information should ideally be concentrated in different subspaces. We describe subspace patching in detail in App. D.1 and address potential concerns from (Makelov et al., 2023) in App. G.2. In sum, we aim to explore whether the high-level “variable” lies in the same subspace no matter what value it takes by measuring inequality or concentration of subspace patching effect.
@@
We construct a test suite, making use of known circuits in GPT2 Small. It consists of 5 tests. The first 4 tests are based on the IOI circuit (Wang et al., 2022), where we intervene on the following pieces of information respectively: (1) previous token information contained in layer 4 residual stream (2) subject position in layer 6 (3) subject position and (4) subject name in layer 8 residual stream. The last test is based on Greater-than circuit (Hanna et al., 2023): (5) information about the last two digits of the starting year in layer 9. We measure patching effect as the change in logit difference for IOI-based tests and change in valid answer probability for the Greater-than-based test. See more details in in App. D.2. We then quantify the inequality of patching effect using Gini coefficient. Though very unconventional in machine learning, we find Gini coefficient a suitable metric. Because we want to measure how strongly patching effect is focused on just one or a few subspaces, this notion is captured by Gini coefficient as it measures inequality or degree of concentration. If the target information is mainly encoded in one subspace, the patching effect for this one would be much higher than others, thus resulting in Gini coefficient close to 1. On the other hand, evenly distributed patching effect results in 0. In App. D.4 we show examples of patching effect distributions and their corresponding Gini coefficients, from which we conclude that the coefficient should be >0.6>0.6 to be considered satisfactory (i.e., the effect is mainly concentrated in one subspace). We also calculate the Gini coefficient over patching effect normalized by dimension $d_{s}$ and variance $\text{Var}_{s}$ of subspaces.
@@
- $d_{s}$ $\text{Var}_{s}$ - $d_{s}$ $\text{Var}_{s}$ - $d_{s}$ $\text{Var}_{s}$ - $d_{s}$ $\text{Var}_{s}$ - $d_{s}$ $\text{Var}_{s}$
@@
Table 1: Results on GPT-2 Small test suite, higher is better. Test 1-5 measure the concentration of various target information in various places in GPT-2 Small (details in App. D.2) based on ground truth circuits (Wang et al., 2022; Hanna et al., 2023). Numbers are Gini coefficient over subspace patching effects (“-”), over effects normalized by subspace dimension ($d_{s}$”), and over effects normalized by subspace variance ($\text{Var}_{s}$”). Last column is the average values over the whole row. We compare NDM with 5 baselines. An identity matrix (“Identity”); a random orthogonal matrix (“Random”); PCA 1 and 2 are matrices composed of eigenvectors after applying PCA on activations. We use dimension configuration $d_{1},\dots,d_{S}]$ of NDM result as the configuration for all baselines. $d_{1},\dots,d_{S}]$ is sorted in descending order, eigenvectors can be sorted in either way according to eigenvalues, resulting in PCA 1 and 2 – meaning either that the smallest subspace receives the biggest (“PCA 1”) or smallest (“PCA 2”) directions of variance; and subspaces obtained by clustering SAE features (“Feature”), using a method adapted from (Engels et al., 2024)) (details are in App. F.2). Importantly, as shown in App. D.4, the degree of concentration of target information becomes satisfactory when Gini coefficient >0.6>0.6.